# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v6

@@ 5 Credit System in Local Economy
Notes: This table summarizes the statistics on the sales in the $326$ retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. The number of stores reported the share of regular customers (column 2) are $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel (firewood and charcoal) categories, respectively. The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual total sales in 1935. The definition of the social classes (“Upper”; “Middle”; “Other”) is not exactly defined in the report. The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935. The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius. The department stores and retail markets are not included in this survey. Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
@@
The information on the credit purchase is considerably personal, as in the case of state-contingent loans in the rural credit market (Udry 1994), and can hardly be documented. Therefore, systematic information on credit purchases in the prewar period is unavailable. Fortunately, however, a survey report on retail shops in 1935 published by the Tokyo City Chamber of Commerce and Industry remained. Table 9 summarizes the information on the sales in those retail shops of rice, fish, fruit and vegetables, and fuels. [^70] This survey covers 939 shops and was conducted on 31st December 1935. I digitized the statistics for all the shops selling the representative items (rice, fish, fruit and vegetables, and fuels) in the credit purchases. The survey instrument includes detailed questions regarding sales and purchases, making it the most comprehensive of the available documents (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by the scale of businesses. I confirm that the main findings are materially similar. First, many annual sales are from their regular customers, most of whom are from the non-upper classes. For example, $73$% of annual sales in rice retailers are from their regular customers, and more than $80$% of them are from non-upper classes, including the factory workers. Second, and most importantly, credit purchase is the popular sales method, especially among rice and fuel retailers: more than $70$% of the sales were done with credit in both shops. Although cash was also used frequently among greengrocers, $34$% of the sales were settled using credit. This supports the evidence that credit purchase was a popular payment method, and regular customers settled a large part of it. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions in the context of a long-term relationship with their regular customers. The final column of this table further provides suggestive evidence behind this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (roughly $330$ meters) radius. As shown, rice and fuel stores have 2-3 more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
by Different Time-series Frequencies
@@
Note: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively. Figure 6b and 6d illustrate the time-series plots of the average daily and semi-monthly credit purchase minus payment. Figures 6a and 6b show the daily data between 12th January and 11th November 1919. Figures 6c and 6d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919. All figures are based on the 30 THBS households that used any credit purchases. Source: Created by the author using the THBS sample.
@@
Next, I look into consumers’ credit purchase behavior using the THBS dataset. Most of the THBS households ($29/33$) used any credit purchases during their sample period, supporting the evidence that this was a common practice among factory workers in Tsukishima. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the households’ behavior of making the payments for the credit purchase on or soon after their paydays. The purchases series shows a similar tendency but with lower spikes and larger variance. This indicates that the households purchased goods by a certain amount around their paydays and then used it when they needed unconservable food such as fish, fruit, and vegetables. Figure 6b illustrating the net purchases confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figures 6a and 6d show that the average payment exceeds the average purchase in almost all semi-month cells, meaning that many households were able to repay their loans. This is consistent with the findings from the survey report that credit purchases function based on the long-term relationship between retailers and consumers in the local economy.
@@
Number of THBS households 33 33 33 33 34 33
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets. Notes: The dependent variable in Columns (1)-(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users). The dependent variable in Columns (4)-(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases. Columns (1)-(2) and (4)-(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model. “CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings. The sample mean of the CV is $5.76$ (Std. Dev. = $5.72$). All the regressions include the family composition variables: average family size, share of children aged 6–12, share of children aged 13–16, and share of adults aged 17+ (share of children aged 0–5 is used as a reference group). The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6. $F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)-(2) and (4)-(5). $\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6). The standard errors for the LPM are based on the HC2 estimator proposed by Horn et al. (1975).
@@
Column (1) of Table 10 presents the result for equation 4. The estimated coefficient is close to zero and statistically insignificant. A potential concern herein may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households using a relatively smaller amount of credit purchases, say less than the first quartile of the average credit purchases. Columns (4)-(5) summarize the results in the same column layouts as Columns (1)-(3), confirming that the estimates are still close to zero in all the specifications. Finally, the final row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households to use credit purchases. This suggests that retailers may not have put weight on the volatility of earnings when trading with factory workers on credit. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit purchases.
@@
The results so far indicate that retailers have been generous in the earnings volatility risk. This, in turn, implies that retailers may factor risk in the form of interest into their loans or retail prices. Unfortunately, there are no systematic statistics on the interest in credit purchases. Thus, I calculate the risk premiums backward by differencing payments and purchases. This backward calculation technique requires information both on the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were sold in greengrocers (yaoya) in many cases but sometimes sold in variety stores (aramonoya). Fortunately, I found that a specific month in a budget book has documented the price-by-retailer information. [^72] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and variety stores between 1st and 14th February 1919 and paid exactly the same amounts on the 15th. In the latter half of February, the household paid the same amounts to greengrocers and variety stores on 28th, but less amount to the fish stores. Since the budget book for March does not have information on the type of retailers, it is difficult to know exactly how the reimbursement shortfall was handled. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
To summarize, the available documents show that retailers competed with other retailers in the same sales area. Under this condition, they provided regular consumers with the opportunity to trade on credit without interest for at least one month. This ease of accessibility to credit purchases might enable workers to smooth consumption in response to adverse idiosyncratic shocks in the short-run.
@@ 6 Conclusion
The estimated elasticity for total expenditure is in a similar range to those in rural economies in developing countries, although only a few available references exist. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was roughly $0.5$. As for the urban households, the estimated elasticity is $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, one must be careful when making comparisons with these estimates based on the urban samples, including a variety of occupations. At least, however, the factory workers in the historical metropolis of Japan around WWI had been relatively vulnerable to the idiosyncratic income shocks.
@@
The finding on the attenuation bias due to the measurement errors is potentially important in that they may prompt a reinterpretation of existing results. Errors in the timing of payroll can result in inconsistencies in the aggregate data for income and expenditure, albeit the household-level studies often use aggregate data. In this case, modifying the aggregation period to align with wage payments can reduce attenuation bias. Even if the aggregation period is unadjustable, clarifying payroll timing and discussing potential bias is useful.
@@ A.1 Developments of the Cities
Note: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities. Figure A.1b shows the share of people in Tokyo (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan. Sources: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
@@
Table A.1: City Characteristics:Tokyo, Osaka, Kyoto, and Nagoya
@@
Notes: 1. Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed. 2. Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation. 3. Panel C: The number of electrical lights measured at the end of 1921 is divided by the number of people in the 1920 Population Census. The number of houses accessible to water supply measured at the end of September 1920 is divided by the number of total households measured in the 1920 Population Census (%). 4. Panel D: The average daily temperature and precipitation between 1st January 1918 and 31st December 1923 and between 1st February 1920 and 30th April 1920 are listed. Sources: The number of people in 1920 is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of male workers and their industrial structures are from the Statistics Bureau of the Cabinet (1929a, pp. 20–21; 26–29). The number of houses accessible to the water supply is from the Water Works Association (1922, pp. 69–70). The total number of households is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of electrical lights is from the Ministry of Posts and Telecommunications (1922, p. 558) The daily temperature and precipitation are from the Japan Meteorological Agency (database).

### v7

@@ 5 Credit System in Local Economy
Notes: This table summarizes the statistics on the sales in the $326$ retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. The number of stores reported the share of regular customers (column 2) are $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel (firewood and charcoal) categories, respectively. The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual total sales in 1935. The definition of the social classes (“Upper”; “Middle”; “Other”) is not exactly defined in the report. The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935. The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius. The department stores and retail markets are not included in this survey.
Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
@@
The information on the credit purchase is considerably personal, as in the case of state-contingent loans in the rural credit market (Udry 1994), and can hardly be documented. Therefore, systematic information on credit purchases in the prewar period is unavailable. Fortunately, however, a few official survey reports documented some information on this payment method. A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers. [^75] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1). Those include not only retailers but also wholesale and other small businesses, and the number of respondents varies depending on the question items. The number of respondents in the statistical table on the payment methods is $2,654$, of which $1,197$ are retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo around 1930. For example, 99%(155/157) of grain retailers, 81%(35/43) of fish retailers, 78%(64/82) of greengrocers, and 94%(45/48) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp. 260; 263). A survey report on retail shops in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information. Table 9 summarizes the information on the sales in those retail shops of rice, fish, fruit and vegetables, and fuels. [^76] This survey covers 939 shops and was conducted on 31st December 1935. I digitized the statistics for all the shops selling the representative items (rice, fish, fruit and vegetables, and fuels) in the credit purchases. The survey instrument includes detailed questions regarding sales and purchases, making it the most comprehensive of the available documents (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by the scale of businesses. I confirm that the main findings are materially similar. First, many annual sales are from their regular customers, most of whom are from the non-upper classes. For example, $73$% of annual sales in rice retailers are from their regular customers, and more than $80$% of them are from non-upper classes, including the factory workers. Second, credit purchase is a popular sales method, especially among rice and fuel retailers: more than $70$% of the sales were done with credit in both shops. [^77] These figures are systematically lower than those from the 1930 Survey Report. Given that both surveys show a similar tendency regarding the use of credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself. Although cash was also used frequently among greengrocers, $34$% of the sales were settled using credit. This supports the evidence that credit purchase was a popular payment method, and regular customers settled a large part of it. [^78] The 1930 Survey reveals that the retailers had also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157). This means that credit purchase was a familiar payment method for retailers as well as consumers. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions in the context of a long-term relationship with their regular customers. [^79] A textbook for retailers published in 1924 explains an interesting mechanism behind this long-term relationship from an aspect of customer psychology: credit purchase is made not only for the convenience of the customer’s budget but also because of the “vanity” of being able to be a regular customer (Shimizu 1924, pp. 115–116). The final column of this table further provides suggestive evidence behind this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (roughly $330$ meters) radius. As shown, rice and fuel stores have 2-3 more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
@@
Note: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively. The blue and red dashed lines indicate the average credit purchases and payments, respectively. Figure 6b and 6d illustrate the time-series plots of the average daily and semi-monthly credit purchase minus payment. The red solid line highlights the zero value. Figures 6a and 6b show the daily data between 12th January and 11th November 1919. Figures 6c and 6d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919. All figures are based on the 29 THBS households that used any credit purchases.
Source: Created by the author using the THBS sample.
Next, I look into consumers’ credit purchase behavior using the THBS dataset. Most of the THBS households ($29/33$) used any credit purchases during their sample period, supporting the evidence that this was a common practice among factory workers in Tsukishima. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the households’ behavior of making the payments for the credit purchase on or soon after their paydays. The purchases series shows a similar tendency but with lower spikes and larger variance. This indicates that the households purchased goods by a certain amount around their paydays and then used it when they needed unconservable food such as fish, fruit, and vegetables. Figure 6b illustrating the net purchases confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figure 6c shows that the payment slightly exceeds purchase, suggesting that many households could more or less repay their loans. Although the difference between payment and purchase is relatively obvious after August 1919, Figure 6d indicates that this difference is not statistically significant in all the semi-months. [^80] Strictly, payment can include accrued liabilities that should have been paid prior to the previous period. It is, therefore, reasonable that the average payment is greater than the average purchase in most semi-month cells. Thus, this is not a systematic trend but may be due to the smaller available number of households in the later sample period (Online Appendix Figure B.4).
@@
Number of THBS households 33 33 33 33 33 33
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets.
@@
Notes: The dependent variable in Columns (1)-(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users). The dependent variable in Columns (4)-(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases. Columns (1)-(2) and (4)-(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model. “CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings. The sample mean of the CV is $5.76$ (Std. Dev. = $5.72$). All the regressions include the family composition variables: average family size, share of children aged 6–12, share of children aged 13–16, and share of adults aged 17+ (share of children aged 0–5 is used as a reference group). The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6. $F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)-(2) and (4)-(5). $\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6). The standard errors for the LPM are based on the HC2 estimator proposed by Horn et al. (1975).
@@
Column (1) of Table 10 presents the result for equation 4. The estimated coefficient is close to zero and statistically insignificant. A potential concern herein may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households using a relatively smaller amount of credit purchases, say less than the first quartile of the average credit purchases. Columns (4)-(5) summarize the results in the same column layouts as Columns (1)-(3), confirming that the estimates are still close to zero in all the specifications. Finally, the final row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households to use credit purchases. This suggests that retailers may not have put weight on the volatility of earnings when trading with factory workers on credit. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit purchases. [^82] Despite this, low-income households might have had more difficulty using credit purchases. A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp. 346–355).
@@
The results so far indicate that retailers have been generous in the earnings volatility risk. This, in turn, implies that retailers may factor risk in the form of interest into their loans or retail prices. Unfortunately, there are no systematic statistics on the interest in credit purchases. Thus, I calculate the risk premiums backward by differencing payments and purchases. This backward calculation technique requires information both on the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were sold in greengrocers (yaoya) in many cases but sometimes sold in variety stores (aramonoya). Fortunately, I found that a specific month in a budget book has documented the price-by-retailer information. [^83] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and variety stores between 1st and 14th February 1919 and paid exactly the same amounts on the 15th. [^84] While the retailers used credit purchases from wholesalers (footnote 78), the repayment period was often set around one month (Tokyo City Office 1932, pp. 157-159). Therefore, this makes sense for consumers to make repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and variety stores on 28th, but less amount to the fish stores. Since the budget book for March does not have information on the type of retailers, it is difficult to know exactly how the reimbursement shortfall was handled. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
Finally, I assess the loss for the retailers in credit purchases. According to the statistics on the losses on credit sales documented in the 1930 Survey Report, almost all of the retailers had suffered losses in credit purchases. For example, 99% of rice and 97% of fuel retailers more or less lost their sales. This means that credit purchase was inherently a loss-making sales method for the retailers. Since systematic statistics are unavailable to clarify the percentage of total profits corresponding to these losses. Thus, it is impossible to discuss how severe these losses are for the retailers. However, an important fact is that these losses were usually capped by 5% of total sales. For example, the available statistics suggest that 87% of rice and 65% of fuel retailers suffered losses amounting to 5% of sales. This supports the evidence that while the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent. This seems to be a plausible sales strategy for retailers to deal with adverse selection issues under imperfect consumer information. Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail.
To summarize, the available documents show that retailers competed with other retailers in the same sales area. Under this condition, they provided regular consumers with the opportunity to trade on credit without interest for at least one month. The 1930 Survey Report indeed stated that “intense competition among competitors has led to a high frequency of credit purchases” (Tokyo City Office 1932b, p. 415). This accessibility to credit purchases could help workers manage consumption after unexpected individual setbacks in the short term. [^85] Another potential explanation for the popularity of credit purchases might be thefts. However, the number of thefts $100$ people in 1919 is $0.5$ in Tsukishima, which is close to or slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245). In fact, Tsukishima Police Department does not mention that Tsukshima was an area with a particularly high rate of thefts (Tsukishima Police Department 1976, pp. 244–245). This suggests that this crime channel is not relevant to explain the use of credit purchases.
@@ 6 Conclusion
The estimated elasticity for total expenditure is in a similar range to those in rural economies in developing countries, although only a few available references exist. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was roughly $0.5$. As for the urban households, the estimated elasticity is $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, one must be careful when making comparisons with these estimates based on urban samples, which include a variety of occupations. At least, however, the factory workers in the historical metropolis of Japan around WWI had been relatively vulnerable to the idiosyncratic income shocks.
@@
The finding on the attenuation bias due to measurement errors is potentially important because it may prompt a reinterpretation of existing results. Errors in the timing of payroll can result in inconsistencies in the aggregate data for income and expenditure, albeit the household-level studies often use aggregate data. In this case, modifying the aggregation period to align with wage payments can reduce attenuation bias. Even if the aggregation period is unadjustable, clarifying payroll timing and discussing potential bias is useful.
@@ A.1 Developments of the Cities
Note: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities. Figure A.1b shows the share of people in Tokyo (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan.
Sources: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
Table A.1: City Characteristics:
Tokyo, Osaka, Kyoto, and Nagoya
@@
Notes:
1. Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed.
2. Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation.
3. Panel C: The number of electrical lights measured at the end of 1921 is divided by the number of people in the 1920 Population Census. The number of houses accessible to water supply measured at the end of September 1920 is divided by the number of total households measured in the 1920 Population Census (%).
4. Panel D: The average daily temperature and precipitation between 1st January 1918 and 31st December 1923 and between 1st February 1920 and 30th April 1920 are listed.
Sources: The number of people in 1920 is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of male workers and their industrial structures are from the Statistics Bureau of the Cabinet (1929a, pp. 20–21; 26–29). The number of houses accessible to the water supply is from the Water Works Association (1922, pp. 69–70). The total number of households is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of electrical lights is from the Ministry of Posts and Telecommunications (1922, p. 558) The daily temperature and precipitation are from the Japan Meteorological Agency (database).